OpenStudy (anonymous):

Suppose that f(0)=1 and f′(x)≤3 for all values of x. Use the Mean Value Theorem to determine how large f(4) can possibly be f(4)≤___________

OpenStudy (anonymous):

@sourwing help

OpenStudy (anonymous):

well, the mean value theorem says f(b) - f(a) = f(c) (b-a) where a = 0 and b = 4. What can you say?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

f(4) - f(0) = f'(c) (4 - 0), f(4) = f'(c) (4) + f(0) you want to maximize f(4)

OpenStudy (anonymous):

and the answer is?

OpenStudy (anonymous):

you want f'(c) to be as big as possible right? so let f'(c) = 3 f(4) = 3(4) + 1 = 13 so f(4) <= 13

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@sourwing, could you help me in a couple questions more?

OpenStudy (anonymous):

no, but post your question and someone might answer it

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

here is another way to think about this easy at noon you are at mile marker one you can walk no more than 3 mph how far can you be at 4 pm?

OpenStudy (anonymous):

you cannot have walked more than \(3\times 4=12\) miles so you cannot be past mile marker \(13\)